solve by the substitution method 7m+n=10 m-6n=26 the solution is.
-7x+3y=4 -2x+y=10 the solution
9x+8y=29 -7x+y=28 the solution
I really need some help. I dont understand how to do this

Answer by Alan3354(69443)

solve by the substitution method
7m+n=10
m-6n=26 --> m = 6n + 26
Sub for m in the 1st eqn
7m + n = 10
7(6n+26) + n = 10
42n+182 + n = 10
43n = -172
n = -4
-----------
m = 6n + 26
m = 2
============
-7x+3y=4 -2x+y=10
Solve for y in the 2nd eqn, sub into the 1st.
============
9x+8y=29 -7x+y=28
Same thing.